OpenStudy (liliakarina):

In order to transform this straight line equation: y = x + 5, to its polar form with the radius vector 'r' in function of the vectorial theta angle. In order to transform this straight line equation: y = x + 5, to its polar form with the radius vector 'r' in function of the vectorial theta angle. @Mathematics

OpenStudy (chriss):

I am not exactly sure what you are asking, but \[r \sin (\theta)=r \cos (\theta)+5\]

OpenStudy (chriss):

\[y=r \sin (\theta)\] and \[x=r \cos (\theta)\]

OpenStudy (chriss):

I guess you are trying to find \[(r,\theta)\]

OpenStudy (liliakarina):

That's exactly what I was trying to figure out. Sorry 'cause I did not formulate correctly the answer. Thanks.

OpenStudy (chriss):

I'm working through it. I have r. It is \[r=\frac{5}{\sin(\theta)-\cos(\theta)}\] I arrived at it like this: \[rsin(\theta)=rcos(\theta)+5\] \[rsin(\theta)-rcos(\theta)=5\] \[r(\sin(\theta)-\cos(\theta))=5\] \[r=\frac{5}{\sin(\theta)-\cos(\theta)}\] Give me a moment on theta.

OpenStudy (liliakarina):

Cool, okay.

OpenStudy (chriss):

I'm a little stuck here. I'm not really sure how to isolate theta in this equation. I'll keep at it, but I don't have a solution for theta at the moment.

OpenStudy (liliakarina):

It's okay, thank you.

OpenStudy (chriss):

you're welcome... at the moment I have it down to \[\sin(\theta)-\cos(\theta)=\frac{5}{r}\] I'm just not sure where to go with it from there.

OpenStudy (liliakarina):

All right, I'll figure out from here.